OpenStudy (bezerkhd):

Solve 4(x + 1) = 4x + 4

OpenStudy (anonymous):

Distribution.

OpenStudy (anonymous):

4*x is what? 4* 1 is what?

OpenStudy (xmissalycatx):

Hello there! All you need to do is distribute the 4 to everything in the parentheses! :) So you'd get 4x + 1 = 4x + 1 which is a true statement.

OpenStudy (xmissalycatx):

Must you solve for x?

OpenStudy (bezerkhd):

Sorry about that. There were no notifications.

OpenStudy (dayakar):

for all values of x the equation is true there fore it has infinity solutions

OpenStudy (bezerkhd):

Possible answers: a) No solution b) 0 c) All real numbers d) 1

OpenStudy (lanmanna):

No solution.

OpenStudy (anonymous):

4(x + 1) = 4x + 4 4x + 4 = 4x + 4 (distributive property) 4 = 4 (subtract 4x from both sides) Solution is all real numbers

OpenStudy (lanmanna):

4(x + 1) = 4x + 4 4x+4 = 4x + 4 4x-4x = 4-4 0 = 0